Tim earns twice as much each week as a tutor than he does pumping gas. His total weekly wages are $150 more than that of his sister. She earns one quarter as
Much as Tim does as a tutor. How much does tim earn as a tutor?

Answer by scott8148(6628)

"Tim earns twice as much each week as a tutor than he does pumping gas" ___ t = 2g

"$150 more than that of his sister. She earns one quarter as Much as Tim does as a tutor" ___ t + g = 150 + (t/4)

clearing fraction ___ 4t + 4g = 600 + t

substituting ___ 4t + (2t) = 600 + t

5t = 600 ___ t = 120